I need som help on solving this question please.
2(-5+12i)/1+i

Answer by stanbon(75887)

2(-5+12i)/(1+i)
-------------------------------------
Notice that (1+i)(1-i) = 1-i+i+1 = 2
----------------------------------------------
Multiply numerator and denominator by 1-i to get:
[2(-5+12i)(1+i)/(2]
---
= (5-12)+(12i-5i)
= -7+7i
= -7(1+i)
===============
Cheers,
Stan H.
===============